oreo:

Wastewater is filling barrels at the rate of 13 quarts per hour. The recycling facility picks up 96 full barrels on each trip, and each barrel holds 30 quarts of wastewater. How many days are between each pickup? Show your work to receive full credit.

Mercury:

old question but will respond so this can be closed Assuming the wastewater gets picked up as soon as they reach 96 barrels: \[(\frac{ 96~barrels }{ 1~pickup })(\frac{ 30~quarts }{ 1~barrel })(\frac{ 1~hr }{ 13~quarts })(\frac{ 1~day }{ 24~hrs })\]notice how the units cancel out and you're left with days/1 pickup